Help, I'm very lost - if I could get help solving this one the rest should come back to me, I hope.
6.2(x - 15.1) = 9.3(-x - 3.2)
Thank you for your help.

Found 2 solutions by ankor@dixie-net.com, Genius:
Answer by ankor@dixie-net.com(12701)

6.2(x - 15.1) = 9.3(-x - 3.2)
:
Mult what's inside the brackets first:
6.2x - 93.62 = -9.3x - 29.76
:
If decimals bother you, mult eq by 100 to get rid of them,
620x - 9362 = -930x - 2976
:
Perform the necessary operations to get the variable on the left and the numerical values on the right:
:
Add 9362 to both sides:
620x = -930x - 2976 + 9362
:
Add 930x to both sides, add the numerical values
620x + 930x = + 6386
:
Add the x's
1550x = 6386
:
Divide both sides by 1550
x = 6386/1550
:
x = 4.12
:
:
Check our solution by substituting 4.12 for x in the original equation:
6.2(x - 15.1) = 9.3(-x - 3.2)
:
6.2(4.12 - 15.1`) = 9.3(-4.12 -3.2)
:
Do the math inside the brackets and you have:
6.2(-10.98) = 9.3(-7.32)
:
-68.076 = -68.076 confirms our solution of x = 4.12.
:
A good idea to always check your solution in the original equation.


Answer by Genius(17)